OpenStudy (anonymous):

Which of the statements about the following quadratic equation is true? 6x2 - 8 = 4x2 + 7x The discriminant is greater than zero, so there are two real roots. The discriminant is greater than zero, so there are two complex roots. The discriminant is less than zero, so there are two real roots. The discriminant is less than zero, so there are two complex roots.

OpenStudy (anonymous):

we could write that equation as : 6x² -8 - (4x² + 7x ) = 0 6x² - 8 -4x² -7x =0 2x² -7x - 8 =2 discriminant = b² - 4 ac discriminant = 49 - 4 (2)(-8) discriminant = 49 - ( - 64 ) = 163 The discriminant is greater than zero, so there are two real roots.

OpenStudy (anonymous):

Oh okay. Thanks for ur help.